OpenStudy (anonymous):

Find an equation of the line that has the given slope and passes through the given point: m=5/2,(2,-2)

OpenStudy (anonymous):

Use point slope form: y-y1=m(x-x1) y-(-2)=5/2*(x-2)

OpenStudy (anonymous):

y = 2.5x - 7

OpenStudy (anonymous):

Thank you